A cooking pan whose inner diameter is 20 cm is filled with water and covered with a 0.5 kg lid. If the local atmospheric pressure is 101 kPa, determine the temperature at which the water will start boiling when it is heated.
1
Expert's answer
2012-06-07T08:01:08-0400
R=0.0001
V=0.5/0.0001
P/(V*d)=T=101000/(5000*0.2)=101